
We installed the 3 higher-resolution grisms (B4K, V4K, R4K) during the afternoon in addition to the lower-resolution grisms (B2K and R2K) that are normally installed. When we went to he first object of the night, we ran into a problem where the disperser wheel would time out and get stuck between positions. We moved to zenith so we could enter the Cas cage and reset the wheel, but the problem kept occurring. We then removed the B2K grism to have better balance in the disperser wheel and we observe without further problems. This was somewhat strange as we had the same setup in April and the disperser wheel worked without problems with the 5 grisms installed. It may be a factor that it was 8C warmer in April than now. We may want to see if there is any way to lubricate the disperser wheel. Another solution would be to fabricate counterweights to place on the disperser wheel so it is balanced. This would require design and fabrication time. The counterweights would need to be about 1kg so I don't know if that will be feasible.
There were also some problems centering the guider at the start of the night. Hernan T is going to talk with Manuel M tomorrow. Omar provided some help over the phone.
Overall COSMOS behaved during the night. We took spectra of standard stars with the 4 installed grisms to provide data for the ETC lookup tables.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-05-20 | 0.0 | 10.75 | 0.0 | 0.0 | 0.0 | 10.75 |
Links
[1] mailto:spoints@ctio.noao.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/10390/delete?destination=print/content/enr-ctio-blanco-4-m-2019-05-20
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/10390/edit?destination=print/content/enr-ctio-blanco-4-m-2019-05-20
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/15713?destination=print/content/enr-ctio-blanco-4-m-2019-05-20
None.